OpenStudy (anonymous):

HELP SHOWWORK show work please x/8 - 1/6 = 1/6 and x^2 + 8^2= 10^2 bonus: 3x-5

OpenStudy (anonymous):

\[\left\{\frac{x}{8}-\frac{1}{6}=\frac{1}{6},\frac{x}{8}=\frac{1}{6}+\frac{1}{6},\frac{x}{8}=\frac{1}{3},x=\frac{8}{3}\right\} \]\[x^2+64=100,\{x=6,x=-6\} \]\[3x-5<x+11,x<8 \]